### v1

# Dual effect of cholesterol on interfacial water dynamics in lipid membranes: Interplay between membrane packing and hydration

Abstract
Water contained within biological membranes plays a critical role in maintaining the separation between intracellular and extracellular environments and facilitating biochemical processes. Variations in membrane composition and temperature lead to phase state changes in lipid membranes, which in turn influence the structure and dynamics of the associated interfacial water. In this study, molecular dynamics simulations were performed on binary membranes composed of dipalmitoylphosphatidylcholine (DPPC) or palmitoyl sphingomyelin (PSM) mixed with cholesterol (Chol). To elucidate the effects of Chol on interfacial water, we examined the orientation and hydrogen-bonding behavior of water molecules spanning from the membrane interior to the interface. As the Chol concentration increased, a transient slow down in water dynamics was observed in the gel phase at 303 K. Conversely, at higher Chol concentrations, water dynamics were accelerated relative to pure lipid membranes across all temperatures studied. Specifically, at a Chol concentration of 50%, the hydrogen bond lifetime in DPPC membranes decreased to approximately 0.5-0.7 times that of pure lipid membranes. This nonmonotonic behavior is attributed to the combined effects of membrane packing induced by Chol and a reduced density of lipid molecules in the hydrophilic region, offering key insights for modulating the dynamical properties of interfacial water.
I Introduction
The lipid bilayer, a fundamental component of biological membranes, is primarily composed of amphiphilic lipid molecules. These molecules self-assemble through intermolecular interactions into a stable bilayer structure in aqueous environments, with hydrophilic headgroups facing the surrounding water and hydrophobic tails oriented toward the membrane interior. The physical properties of lipid membranes are highly sensitive to both their composition and temperature.[1, 2] Furthermore, the structure and dynamic characteristics of the interfacial water, namely, water molecules interacting with lipid headgroups, varies with the membrane’s phase state, thereby affecting associated biochemical processes. [3, 4]
Cholesterol (Chol) is a representative steroid molecule present of animal cell membranes, accounting for approximately 20% to 50% of their lipid content. Its rigid ring structure and polar hydroxyl group allow Chol to integrate stably into lipid bilayers, where it promotes ordering of the hydrophobic acyl chains, leading to increased membrane condensation. [5, 6, 7, 8] These structural effects of Chol on the membrane also modulate the properties of interfacial water. [9, 10]
Experimentally, various techniques have been employed to investigate the behavior of interfacial water around lipid membranes, including fluorescence spectroscopy, Overhauser dynamic nuclear polarization (ODNP) NMR, quasi-elastic neutron scattering, and sum frequency generation (SFG) spectroscopy. [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] In particular, Chol has been reported to markedly influence the behavior of intefacial water. Cheng et al. reported, based on ODNP measurements, a reduction in water diffusivity within the hydrophobic region of membranes, along with an increase in surface diffusivity under high Chol concentrations. [13] Furthermore, Orlikowska-Rzeźnik et al. investigated the effect of Chol on the hydration of membranes using fluorescence spectroscopy with Laurdan probes and heterodyne-detected vibrational SFG spectroscopy. Their findings indicated that the incorporation of Chol reduces water content within the membrane interior while increasing the orientational ordering of water molecules around the lipid headgroups. [22, 23]
Numerous molecular dynamics (MD) simulations have been conducted to investigate the lipid membrane–water interface. [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 7, 35, 36, 37, 38, 39, 40] MD simulation studies have also examined how Chol influences the structural organization of lipid membranes, including its effects on bilayer thickness, lipid packing, and phase behavior. [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] Specifically, clarifing the influence of Chol on the hydrogen-bond (H-bond) network within the surface hydration layer is essential for understanding membrane–water interactions at the microscopic level. Elola et al. reported that increasing the Chol concentration up to 50% enhances the diffusivity of water molecules near the membrane surface. [49] This enhancement in water mobility was attributed to the disruption of lipid-water-lipid H-bonds, while promoting the formation of water-water H-bond network. Oh et al. characterized the structure of interfacial water by modeling the HB network at lipid membrane surfaces containing Chol using a graph-based approach. [53] Their analysis indicated that increasing the Chol concentration enhances the connectivity of the H-bond network among water molecules, resulting in a structure that more closely resembles bulk-like behavior.
We previously performed MD simulations of dipalmitoylphosphatidylcholine (DPPC) and palmitoyl sphingomyelin (PSM) membranes containing 10% Chol as well as pure membranes to elucidate the dynamics of water molecules H-bonded with lipid molecules. [55] While DPPC and PSM share an identical hydrophobic tail structure, they differ in the architecture of their hydrophilic headgroups, particularly in the configuration of their polar functional groups [see Fig. 1(a) and (b)]. Notably, our MD simulations of interfacial water dynamics in two types of lipid bilayers, DPPC and PSM, revealed that water exhibits slower dynamics near PSM compared to DPPC. This behavior is primarily attributed to H-bonds formed between water hydrogen atoms and the oxygen atoms of PSM, labeled $\mathrm{O}^{\mathrm{S4}}$ and $\mathrm{O}^{\mathrm{S5}}$ [see Fig. 1(b)]. A similar retardation of interface water dynamics near PSM membranes has also been reported in previous studies. [56, 57] Furthermore, we found that the Chol slows the interfacial water dynamics, with a more pronounced effect in DPPC than in PSM. In particular, at 303 K, corresponding to the gel phase, the H-bond lifetime between lipid oxygen atoms and water hydrogen atoms increased at a Chol concentration of 10%.
Our findings and those previously reported in MD simulations indicate that Chol exhibits a dual effect of retarding and accelerating the dynamics of water molecules near the membrane interface with a significant dependence of its concentration. In this paper, we elucidate how Chol affects the H-bond dynamics of interfacial water, particularly as a function of Chol concentration.
II MD simulations
Figure 1: Structures of DPPC (a), PSM (b), and Chol (c) molecules studied in this paper.
Table 1: Numbers of lipid (DPPC or PSM), Chol, and water molecules in mixture and pure membrane systems.
$x_{\mathrm{Chol}}$ 0 0.1 0.5
DPPC / PSM 200 200 200
Chol 0 22 200
Water 20000 22200 40000
The structures of the lipid molecules, DPPC, PSM, and Chol are depicted in Fig. 1. The mole fraction of Chol in the lipid membrane was set to $x_{\mathrm{Chol}}=0$, 0.1, and 0.5, with the number of water molecules maintained at a ratio of 100 per total number of lipid and cholesterol molecules (see Table 1). The CHARMM36 force field was employed for the lipid molecules (DPPC, PSM, and Chol), [58] and water was modeled using the TIP3P model optimized for compatibility with the CHARMM force field. [59] All the MD simulations were performed using GROMACS 2022.4. [60] The initial configurations of the lipid membrane systems were constructed using the CHARMM-GUI. [61, 62, 63, 64] System equilibration was carried out following the CHARMM-GUI protocol, as described in our previous study. [55] Three initial configurations were prepared for each $x_{\mathrm{Chol}}$. The temperatures investigated were 303 K and 323 K. The production runs of 10 ns in the $NPT$ ensemble. Each production run was performed three times per configuration with different initial atomic velocities assigned according to the Maxwell–Boltzmann distribution, resulting in a total of nine trajectories for each combination of $x_{\mathrm{Chol}}$ and temperature.
III Results and discussion
III.1 Area per lipid
Figure 2: Snapshots of DPPC-Chol systems at 303 K, with DPPC shown in light blue, cholesterol in violet, phosphorus atoms in orange, and water in transparent blue. The upper panels present top-down views in the $xy$-plane with water omitted, while the lower panels show side views along the $z$-axis. All images were generated using the molecular visualization software VMD. [65] Panels correspond to (a) $x_{\mathrm{Chol}}=0$, (b) $x_{\mathrm{Chol}}=0.1$, and (c) $x_{\mathrm{Chol}}=0.5$.
Figure 3: Area per lipid as a function of the molar fraction of Chol, $x_{\mathrm{Chol}}$, for (a) DPPC and (b) PSM. Circles (blue) and squares (pink) represent the results at 303 K and 323 K, respectively.
Figure 2 presents snapshots of DPPC-Chol systems at 303 K, illustrating the effect of varying Chol concentration, $x_{\mathrm{Chol}}$. In this study, the total number of lipid molecules was fixed, as described in Table 1. Thus, increasing $x_{\mathrm{Chol}}$ led to a corresponding increase in the simulation box size. It is also observed that $x_{\mathrm{Chol}}$ influences the bending behavior of membranes and the orientation of the hydrophobic group of the lipid molecules. The top-down snapshots in Fig. 2 reveal that at $x_{\mathrm{Chol}}=0.1$, Chol is stably incorporated into the membrane and exhibits strong lateral aggregation, with minimal exposure at the membrane interface. In contrast, at $x_{\mathrm{Chol}}=0.5$, the presence of Chol becomes apparent at the membrane interface, which is regarded as the effective dilution of lipid molecules compared to the system with $x_{\mathrm{Chol}}=0.1$.
We examined how the membrane area in the $xy$ plane varies with the Chol concentration. Specifically, we calculated area per lipid, defined as the average area occupied by lipid and Chol molecules projected onto the $xy$-plane. This metric serves as an indicator of the lipid packing density within the membrane. [1] Figure 3(a) and (b) present the area per lipid as a function of $x_{\mathrm{Chol}}$ for the DPPC and PSM systems, respectively. This figure reveals that area per lipid gradually decreases with increasing $x_{\mathrm{Chol}}$, consistent with the results reported in previous MD studies. [41, 43, 44, 7, 46, 39] The reduction in area per lipid indicates that Chol decreases the free volume within the lipid membrane, leading to the condensation effect. Compared to DPPC, the area per lipid in PSM decreases more gradually with increasing $x_{\mathrm{Chol}}$, reflecting the fact that PSM already exhibits a smaller area in the Chol-free system. The difference in the reduction of area per lipid between DPPC and PSM can be attributed to the presence of an oxygen atom in the hydroxyl group of PSM, denoted as $\mathrm{O}^{\mathrm{S4}}$ in Fig. 1(b). In fact, the hydroxyl group of PSM acts as both the H-bond donor and acceptor, [66, 67] contributing to stronger condensation effect through enhanced H-bonding between lipid molecules, particularly for low $x_{\mathrm{Chol}}$. Finally, at $x_{\mathrm{Chol}}=0.5$, area per lipid approaches a saturated value that is nearly identical for both DPPC and PSM, regardless of temperature.
III.2 Water distribution near membrane interface
Figure 4: (a) and (b): Ratios of the water molecule density distribution, $\rho(z^{\prime})$, to the bulk water density, $\rho_{\mathrm{bulk}}=33.50$ and 32.95 $\mathrm{nm}^{-3}$, for DPPC and PSM at 303 K, respectively. (c) and (d): Average numbers of H-bonds, $N_{\mathrm{HB}}$, as a function of $z^{\prime}$, using the same axis as in (a) and (b), at 303 K. (e) and (f): $N_{\mathrm{HB}}$, normalized by the number of nearest neighbor oxygen atoms $N_{\mathrm{NN}}$, as a function of $z^{\prime}$, using the same axis as in (a) and (b), at 303 K. The vertical dashed lines represent the boundaries separating the three regions.
In lipid membranes, water molecules are largely excluded from the membrane interior and predominantly resides near the interfacial region. To examine the spatial distribution of water molecules, it is common to calculate the number density $\rho(z)$ along the $z$-axis, centered at the midpoint of the membrane. However, accurately characterizing interfacial water along the $z$-axis is difficult due to significant fluctuations of the membrane interface using the $z$-axis. Alternatively, a method was proposed to calculate the number density profile of water relative to the lipid headgroups using Voronoi tassellation based on the positions of lipid phosphorus atoms. [31, 33] Furthermore, we developed a simplified algorithm that does not rely on the Voronoi tassellation by defining $z^{\prime}$ as the $z$-axis distance between water molecule and the nearest phosphorus atom projected onto the $xy$-plane. [55] Thus, $\rho(z^{\prime})$ represents the distribution of water molecules while accounting for fluctuations in the membrane interface position. When $z^{\prime}\approx 0$, water molecules are located near the lipid phosphorus atoms. For $z^{\prime}<0$, water molecules is penetrated into the membrane interior, whereas for sufficiently large $z^{\prime}$, the number density $\rho(z^{\prime})$ approaches the bulk water number density, $\rho_{\mathrm{bulk}}$.
Figure 4(a) and (b) show the ratio $\rho(z^{\prime})/\rho_{\mathrm{bulk}}$ at 303 K for DPPC and PSM, respectively. The results of $x_{\mathrm{Chol}}=0$ and 0.1 are consistent with those reported in our previous study. [55] Figure S1 of the supplementary material shows the results at 323 K. Here, the $z^{\prime}$ axis is divided into three distinct regions: the membrane interior (Region 1), the interface (Region 2), and the bulk aqueous phase (Region 3). [31, 33] We defined Region 1 as $z^{\prime}<-0.25$ nm, Region 2 as $-0.25~\mathrm{nm}<z^{\prime}<0.85\mathrm{nm}$, and Region 3 as $z^{\prime}>0.85~\mathrm{nm}$, as indicated by the dashed vertical lines in Fig. 4.
Both DPPC and PSM exhibit a non-monotonic dependence on $x_{\mathrm{Chol}}$ in Region 2 at 303 K. In particular, for DPPC, the number density of water at the membrane interface decreased from $x_{\mathrm{Chol}}=0$ to $x_{\mathrm{Chol}}=0.1$, but increases markedly at $x_{\mathrm{Chol}}=0.5$. This increase in $\rho(z^{\prime})$ is consistent with the results performed by Elola et al. [49] As illustrated in Fig. S1 of the supplementary material, at 323 K, the number density of water in Region 2 shows little variation between $x_{\mathrm{Chol}}=0$ and $x_{\mathrm{Chol}}=0.1$. However, at $x_{\mathrm{Chol}}=0.5$, it increases noticeably, similar to the trend observed at 303 K. This non-monotonic behavior of $\rho(z^{\prime})$ appears to contrast with the results in Fig. 3, which indicate that increasing $x_{\mathrm{Chol}}$ leads to a more condensed lipid membrane state. This apparent discrepancy can be interpreted as follows: Chol, being shorter than typical lipid molecules, resides near the membrane center. While its hydroxyl group and the lipid phosphate groups are in close proximity, they do not extend fully to the membrane interface, particularly at a low Chol concentration. However, at high Chol concentrations, the lateral spacing between lipid headgroups increases, leading to greater exposure of Chol at the interface, which can be observed in the snapshot in Fig. 2. This increased exposure facilitates the accumulation of water molecules at the interface.
Compared to DPPC, PSM exhibits small changes in $\rho(z^{\prime})$ at $x_{\mathrm{Chol}}=0.1$ from $x_{\mathrm{Chol}}=0$, while at $x_{\mathrm{Chol}}=0.5$, it shows a compatible increase in the water number density at Region 2, similar to the behavior observed in DPPC. As mentioned in Sec. III.1, the area per lipid is smaller for PSM than for DPPC in the absence of Chol (see Fig. 3), indicating that PSM forms a more tightly packed membrane. Consequently, the exclusion effect of Chol on water molecules is less pronounced in PSM. Eventually, as the Chol concentration increases, the lateral spacing between lipid headgroups expands, promoting the redistribution of water molecules from the membrane interior toward the interface.
We next examined the H-bond number, $N_{\mathrm{HB}}$, formed by water molecules, as a function of $z^{\prime}$, as is shown in Fig. 4 (c) and (d), for DPPC and PSM, respectively. The corresponding results at 323 K are displayed in Fig. S1 (c) and (d) of the supplementary material. An H-bond was identified based on the geometric criteria: an oxygen-oxygen distance $r_{\mathrm{oo}}<0.35~\mathrm{nm}$ and a hydrogen-donor-acceptor angle $0^{\circ}<\beta<30^{\circ}$[68] Note that when counting $N_{\mathrm{HB}}$, we included not only those between water molecules but also those formed between water molecules and lipid oxygen atoms, ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D6}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, as well as Chol oxygen ($\mathrm{O}^{\mathrm{Chol}}$) (see Fig. 1). The value of $N_{\mathrm{HB}}$ decreases progressively towards the membrane interior in both DPPC and PSM. In Region 1, the presence of Chol noticeably reduces $N_{\mathrm{HB}}$, compared to the Chol-free system.
Figure 5: Two-dimensional plots of joint distribution of $\rho(z^{\prime},\cos\theta)$, normalized by $\rho(z^{\prime})$, at 303K for $x_{\mathrm{Chol}}=0$ [(a) and (d)], 0.1 [(b) and (e)], and 0.5 [(c) and (f)]. Panels (a), (b), and (c) correspond to DPPC, while panels (d), (e), and (f) corresponds to PSM.
Figure 4(c) and (d) show a decrease in the number of H-bonds within the membrane interior (Region 1), which becomes more pronounced in the presence of Chol. This reduction is likely due to the diminished number of water molecules in this region, as indicated in Fig. 4(a) and (b). In contract, in the membrane interface (Region 2), the presence of Chol increases the local water density and is accompanied by a modest increase in the number of H-bonds, particularly at $x_{\mathrm{Chol}}=0.5$. This suggests that Region 2 offers an environment more favorable for the formation of H-bonds between water molecules.
In addition, we analyzed the normalized number of H-bonds formed by oxygen atoms surrounding the water molecules at each position $z^{\prime}$. Specifically, the number of oxygen atoms located within $r_{\mathrm{OO}}<0.35$ nm is denoted as $N_{\mathrm{NN}}$, representing the number of nearest neighbors. The normalized number of H-bonds, $N_{\mathrm{HB}}/N_{\mathrm{NN}}$, was then evaluated as a function of $z^{\prime}$, with the results shown in Fig. 4(e) and (f). Note that $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ denotes the average fraction of neighboring oxygen atoms that form H-bonds at each position $z^{\prime}$[69, 70] In the bulk region (Region 3), the value of $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ is approximately 0.6, indicating that about 60% of the oxygen atoms in the nearest neighbors participate in H-bonding. The value of $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ increases toward the interior of the membrane (Region 1) in the Chol-free system. In other words, when oxygen atoms are present nearby, H-bonds are formed with a higher probability than in the bulk, primarily due to the proximity of oxygen atoms from lipid molecules. However, at $x_{\mathrm{Chol}}=0.5$, the Chol suppressed H-bond formation in the membrane interior for both DPPC and PSM. This suppression can be attributed to the reduced number of H-bonds in that region caused by Chol, as shown in Fig. 4(c) and (d).
III.3 Water orientation near membrane interface
Figure 6: Avarage cosine of the angle $\theta$, $\langle\cos\theta\rangle$, as a function of $z^{\prime}$ for (a) DPPC and (b) PSM, at 303 K. The vertical dashed lines represent the boundaries separating the three regions.
We further analyze how the orientation of water molecules at the membrane interface depends on Chol concentration. Specifically, we calculated the angle $\theta$ between the water dipole vector and the bilayer normal. Previous MD simulations have demonstrated that water orientation varies with the nature of the interacting functional groups. [71, 72, 73, 35, 74, 75, 76] While these results typically represent the $\cos\theta$ dependence on distance $z$ from the membrane center, it should be noted that our analysis uses $z^{\prime}$, a coordinate defined relative to the nearest phosphorus atom. This approach accounts for local fluctuations of the membrane interface and, in particular, enables a more accurate characterization of water orientation near the lipid headgroup region.
Figure 5 shows the joint distribution function $\rho(z^{\prime},\cos\theta)/\rho(z^{\prime})$, which characterizes the orientation of the water dipole vector at position $z^{\prime}$, normalized by $\rho(z^{\prime})$, with varying $x_{\mathrm{Chol}}$ at 303 K. Figure S2 of the supplementary material presents $\rho(z^{\prime},\cos\theta)/\rho(z^{\prime})$ at 323 K. The quantity $\rho(z^{\prime},\cos\theta)/\rho(z^{\prime})$ represents the conditional probability of water molecule orientation at a given position $z^{\prime}$. A value of 0.5 indicates that the probability of a water molecule orienting in a particular direction corresponds to that in the bulk. Values exceeding 0.5 suggest a preferential orientation of water dipoles relative to the bilayer normal at that position.
Figure 5 demonstrates that in Region 3 (bulk region) defined by $z^{\prime}>0.85$ nm, $\rho(z^{\prime},\cos\theta)$ is approximately 0.5 regardless of $\cos\theta$, indicating an isotropic orientation of water molecules. Furthermore, it is demonstrated that both DPPC and PSM exhibit higher values at $-1<\cos\theta<0.5$ in Region 2 defined by $-0.25~\mathrm{nm}<z^{\prime}<0.85~\mathrm{nm}$. This Region 2 likely represents water molecules located near the phosphate groups of the lipid headgroups, as discussed in Sec III.2. The predominantly negative values of $\cos\theta$ suggest that the dipole vectors of these water molecules are oriented towards the membrane interior, consistent with experimental observations. [23] In contrast, within Region 1 defined by $z^{\prime}<-0.25$ nm, water molecules tend to orient upward, and this tendency becomes more pronounced with increasing $x_{\mathrm{Chol}}$.
To highlight how their dipole alignment varies with $x_{\mathrm{Chol}}$, Fig. 6 plots the average orientation $\langle\cos\theta\rangle$ of water molecules as a function $z^{\prime}$ at 303 K. The results at 323 K are illustrated in Fig. S3 of the supplementary material. As mentioned above, in Region 2, water molecules tend to orient their H-atoms toward the lipid membrane; however, the tendency weakens for $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$. This reduction is likely due to the increase in the lateral spacing between lipid headgroups, as discussed in Sec. III.2. Interestingly, at $x_{\mathrm{Chol}}=0.1$, DPPC exhibited a slight increase in water molecule orientation toward the membrane interior, attributed to the condensation effect of Chol, whereas PSM showed almost no change compared to the $x_{\mathrm{Chol}}=0$ case. These results are consistent with the observed $\rho(z^{\prime})$ dependence on $x_{\mathrm{Chol}}$ [see Fig. 4(a) and (b)]. In contrast, within the interior region (Region 1) of the membrane, water molecules tend to orient upward toward the membrane, particularly at $x_{\mathrm{Chol}}=0.5$. This behavior can be attributed to the increased probability of H-bonding, with Chol’s hydroxyl group [denoted $\mathrm{O}^{\mathrm{Chol}}$ in Fig. 1(c)] acting as the acceptor and water molecules as the donors.
III.4 Transition dynamics among three regions
Figure 7: Conditional probability $C_{1,j}(t)$, representing transition dynamics from Region 1 at the initial time $t=0$ to Region 2 until time $t$, or the probability of remaining within the same Region 1 within the time interval $t$ [(a) DPPC at 303 K, (b) PSM at 303 K].
Figure 8: Conditional probability $C_{2,j}(t)$, representing transition dynamics from Region 2 at the initial time $t=0$ to Region 1 or Region 3 until time $t$, or the probability of remaining within the same Region 2 within the time interval $t$ [(a) DPPC at 303 K, (b) PSM at 303 K].
To elucidate the dynamics of water molecules, we analyzed their transition behavior across the three defined regions: interior (Region 1), interface (Region 2), and bulk (Region 3), as shown in Fig. 4(a) and (b). As examined in Ref. 55, we calculated $C_{i,j}(t)$ $(i\neq j)$, the conditional probability that a water molecule initially located in Region $i$ at time 0 reaches Region $j$ by time $t$, having crossed exactly one regional boundary during this interval. $C_{i,i}(t)$ represents the probability that the water molecule remains in Region $i$ without entering the other regions during the time interval between 0 and $t$. Accordingly, the summation over all possible $j$ regions satisfies $\sum_{j}C_{i,j}(t)=1$, ensuring conservation of the total number of water molecules. Note that as $t$ increases, $C_{i,i}(t)$ decreases while $C_{i,j}(t)$ increases.
Figure 7(a) and (b) shows the results of $C_{1,1}(t)$ and $C_{1,2}(t)$ for DPPC and PSM at 303 K, respectively. The corresponding results at 323 K are shown in Fig. S4 of the supplementary material. When $x_{\mathrm{Chol}}$ increases from 0 to 0.1, the water dynamics slowed down, indicating that the time required to transition into Region 2 became longer. This result is consistent with the finding reported in Ref. 55. In contrast, at $x_{\mathrm{Chol}}=0.5$, the transition from Region 1 to Region 2 was accelerated for both DPPC and PSM. As shown in Fig. 4(a) and (b), the number density of water molecules near the boundary between the two regions (at $z^{\prime}=-0.25$ nm) is larger at $x_{\mathrm{Chol}}=0.5$ than at other $x_{\mathrm{Chol}}$ values. This local increase in water density likely facilitates the dynamics of water molecules along the membrane normal. Comparison between DPPC and PSM reveals that the acceleration effect of Chol on water dynamics is particularly pronounced in PSM, as evidenced by $C_{1,1}(t)$ and $C_{1,2}(t)$. Notably, while water dynamics in the Chol-free system, PSM are slower than those in DPPC, the absence of differences in area per lipid and $\rho(z^{\prime})$ between the two systems at $x_{\mathrm{Chol}}=0.5$, as shown in Fig. 3 and Fig. 4, indicates that the local water dynamics become comparable at high Chol concentrations.
Next, we present the results of $C_{2,1}(t)$, $C_{2,2}(t)$, and $C_{2,3}(t)$ for DPPC and PSM at 303 K in Fig. 8(a) and (b), respectively. The corresponding results at 323 K are shown in Fig. S5 of the supplementary material. As time $t$ increases, the value of $C_{2,1}(t)$ saturated in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5 for both DPPC and PSM. Specifically, at $x_{\mathrm{Chol}}=0.5$, $C_{2,1}(t)$ converges to approximately 0.21, which is about twice the value observed at $x_{\mathrm{Chol}}=0$ This result corresponds to $C_{1,2}(t)$ shown in Fig. 7, indicating that the exchange between the membrane interior region (Region 1) and the interface region (Region 2) becomes faster with increasing $x_{\mathrm{Chol}}$. Conversely, the saturated value of $C_{2,3}(t)$ decreased in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5. In particular, at $x_{\mathrm{Chol}}=0.5$, water molecules in Region 2 are more likely to transition into Region 1, thereby relatively reducing the proportion that moves into the bulk region (Region 3). According to $C_{2,1}(t)+C_{2,2}(t)+C_{2,3}(t)=1$, the proportion remaining in Region 2, represented by $C_{2,2}(t)$, decayed slightly more slowly as $x_{\mathrm{Chol}}$ increased.
These results of $C_{i,j}(t)$ indicate that the hydration state in the interface region (Region 2) is significantly influenced by Chol concentration. In particular, the findings suggest that water molecules in Region 2 become more stable at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$ and 0.1.
III.5 H-bond lifetime between water molecules and functional groups
Figure 9: Ratio of H-bond lifetimes, $\tau_{\mathrm{HB}}$, at $x_{\mathrm{Chol}}=0.1$ and 0.50.50.5 to those at $x_{\mathrm{Chol}}=0$, denoted as $\tau_{\mathrm{HB(Chol)}}/\tau_{\mathrm{HB(pure)}}$, for (a) DPPC and (b) PSM.
Finally, to elucidate the interfacial dynamics at a more microscopic level, we analyzed the lifetime of H-bonds formed between water molecules as well as between water molecules and functional groups of DPPC and PSM. The identification of H-bond follows the same criteria described in Sec. III.2. The acceptor oxygen of water molecule is denoted as $\mathrm{O}^{\mathrm{w}}$. Note that the donor oxygen atoms are limited to ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, respectively. This selection is based on the significantly lower number of H-bonds involving other oxygen atoms such as ($\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D6}}$) for DPPC and $\mathrm{O}^{\mathrm{2}}$ for PSM, as confirmed from the joint probability distribution of $r_{\mathrm{OO}}$ and $\beta$ reported in Ref. 55.
We define the H-bond index $h_{i,j}(t)$ is defined as 1 if atoms $i$ and $j$ form a H-bond at time $t$, and 0 otherwise, Using this definition, we calculated the H-bond time correlation function as follows:
$P_{\mathrm{HB}}(t)=\dfrac{\langle h_{i,j}(t)h_{i,j}(0)\rangle}{\langle h_{i,j}(0)\rangle},$ (1)
[…]
We computed $P_{\mathrm{HB}}(t)$ using the Monte Carlo bootstrap method from independent nine trajectories. [80] From the fitting of $P_{\mathrm{HB}}(t)$ to the Kohlrausch–Williams–Watts function, $P_{\mathrm{HB}}(t)=\exp[-(t/\tau_{\mathrm{KWW}})^{\beta_{\mathrm{KWW}}}]$, the H-bond lifetime $\tau_{\mathrm{HB}}$ was evaluated as
[…]
where $\Gamma(\cdots)$ represents the Gamma function.
The values of $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.1$ and 0.5, normalized by the corresponding value at $x_{\mathrm{Chol}}=0$, are plotted as a function of the acceptor oxygen type in Fig. 9. Figure  9 shows that the H-bond lifetime $\tau_{\mathrm{HB}}$ increases at $x_{\mathrm{Chol}}=0.1$, with a particularly pronounced effect in DPPC at 303 K. This is consistent with our previous findings. [55] In contrast, at $x_{\mathrm{Chol}}=0.5$, $\tau_{\mathrm{HB}}$ decreases significantly for both DPPC and PSM. For DPPC, $\tau_{\mathrm{HB}}$ is reduced to approximately 50% at 303 K and 70% at 323 K compared to Chol-free system. Furthermore, the dependence on the type of acceptor oxygen becomes relatively weak. A similar tendency is observed for PSM, with a more pronounced reduction in $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.5$, which exhibits comparable values between 303 K and 323 K.
This acceleration effect of Chol on $\tau_{\mathrm{HB}}$ can be attributed to the increased water content in the membrane interface (Region 2) at $x_{\mathrm{Chol}}=0.5$, as discussed previously [see Fig. 4(a) and (b)]. In addition, the more substantial influence of Chol in PSM is likely due to its intrinsically tight membrane packing in the absence of Chol, compared to DPPC, as noted in Sec. III.1. These findings highlight that Chol concentration modulates the dynamics of H-bonds between water molecules and lipid molecules at the membrane interface in a non-monotonic manner, irrespective of the lipid species.
IV Conclusion
In this study, we investigated how variations in Chol concentration influence the structure and dynamics of interfacial water in DPPC and PSM lipid membranes using MD simulations. Analysis of the area per lipid revealed that increasing Chol concentration, $x_{\mathrm{Chol}}$, leads to a more tightly packed bilayer structure, with this condensation effect being more pronounced in DPPC than in PSM.
Using the distance $z^{\prime}$ between the oxygen atom of a water molecule and the nearby phosphorus atom of lipid, the interfacial water in the lipid membrane was classified into three regions: interior (Region 1), interface (Region 2), and bulk (Region 3). [31, 33, 49, 55] As $x_{\mathrm{Chol}}$ increases, the water number density in the interface region exhibited non-monotonic behavior: it decreased at $x_{\mathrm{Chol}}=0.1$ compared to the Chol-free system, but increased again at $x_{\mathrm{Chol}}=0.5$.
This behavior can be attributed to the dual role of Chol in lipid membranes. While Chol promotes membrane packing, its smaller hydrophilic group compared to lipid headgroups results in dilution at the membrane interface when present in large amounts. As a result, water molecules can more easily access the membrane interface at $x_{\mathrm{Chol}}=0.5$. In other words, although Chol enhances packing and can trap water within the membrane at low concentrations, at higher concentrations, its diluting effect dominates, facilitating water penetration at the interface.
Correspondingly, the dynamics of interfacial water characterized by the transition behavior between the three regions and the H-bond lifetime showed the decrease at $x_{\mathrm{Chol}}=0.1$, followed by an acceleration at $x_{\mathrm{Chol}}=0.5$, consistent with previous MD simulations [49, 53] and spectroscopic experiments [13, 22, 23]. In summary, the non-monotonic behavior in dynamics of water molecules near lipid membrane interface arises from the interplay between membrane packing and increased hydration induced by high Chol concentrations.
SUPPLEMENTARY MATERIAL
The supplementary material presents $\rho(z^{\prime})/\rho_{\mathrm{bulk}}$ and the number of H-bonds, and $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ as a function of $z^{\prime}$ at 323 K (Fig. S1); $\rho(z^{\prime},\cos\theta)$ normalized by $\rho(z^{\prime})$ at 323K (Fig. S2); $\langle\cos\theta\rangle$ as a function of $z^{\prime}$ at 323 K (Fig. S3); $C_{1,j}(t)$ at 323 K (Fig. S4); and $C_{2,j}(t)$ at 323 K (Fig. S5).
Acknowledgements.
This work was supported by JSPS KAKENHI Grant-in-Aid Grant Nos. JP25KJ1764,  JP25K00968, JP24H01719, JP22H04542, JP22K03550, JP21H05249, JP23K26617, JP25K17896, JP24K21230, and JP23H02622. We acknowledge support from the Fugaku Supercomputing Project (Nos. JPMXP1020230325 and JPMXP1020230327) and the Data-Driven Material Research Project (No. JPMXP1122714694) from the Ministry of Education, Culture, Sports, Science, and Technology, the Core Research for Evolutional Science and Technology (CREST) from Japan Science and Technology Agency (JST) (No. JPMJCR22E3), and by Maruho Collaborative Project for Theoretical Pharmaceutics. The numerical calculations were performed at Research Center of Computational Science, Okazaki Research Facilities, National Institutes of Natural Sciences (Projects: 25-IMS-C052).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Data availability statement
The data that support the findings of this study are available from the corresponding author upon request.

### v2

@@
 To elucidate the effects of Chol on interfacial water, we examined the orientation and hydrogen-bonding behavior of water molecules spanning from the membrane interior to the interface.
-As the Chol concentration increased, a transient slow down in water dynamics was observed in the gel phase at 303 K.
+As the Chol concentration increased, a transient slow down in water dynamics was observed in the ripple phase at 303 K.
 Conversely, at higher Chol concentrations, water dynamics were accelerated relative to pure lipid membranes across all temperatures studied.
@@
 Cholesterol (Chol) is a representative steroid molecule present of animal cell membranes, accounting for approximately 20% to 50% of their lipid content.
-Its rigid ring structure and polar hydroxyl group allow Chol to integrate stably into lipid bilayers, where it promotes ordering of the hydrophobic acyl chains, leading to increased membrane condensation.
+For instance, the membranes of human blood cells contain palmitoyl sphingomyelin (PSM) and analogs of dipalmitoylphosphatidylcholine (DPPC), and their Chol content is approximately 50%.
+[5] Owing to its rigid ring structure and polar hydroxyl group, Chol integrates stably into lipid bilayers, where it promotes ordering of the hydrophobic acyl chains, leading to increased membrane condensation.
 [6, 7, 8, 9] These structural effects of Chol on the membrane also modulate the properties of interfacial water.
@@
 [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] Specifically, clarifing the influence of Chol on the hydrogen-bond (H-bond) network within the surface hydration layer is essential for understanding membrane–water interactions at the microscopic level.
-Elola et al.
-reported that increasing the Chol concentration up to 50% enhances the diffusivity of water molecules near the membrane surface.
+Elola and Rodriguez reported that increasing the Chol concentration of 30% and 50% enhances the diffusivity of water molecules near the membrane surface.
 [50] This enhancement in water mobility was attributed to the disruption of lipid-water-lipid H-bonds, while promoting the formation of water-water H-bond network.
@@
 [57, 58] Furthermore, we found that the Chol slows the interfacial water dynamics, with a more pronounced effect in DPPC than in PSM.
-In particular, at 303 K, corresponding to the gel phase, the H-bond lifetime between lipid oxygen atoms and water hydrogen atoms increased at a Chol concentration of 10%.
 Our findings and those previously reported in MD simulations indicate that Chol exhibits a dual effect of retarding and accelerating the dynamics of water molecules near the membrane interface with a significant dependence of its concentration.
@@
 Water 20000 22200 40000
-The structures of the lipid molecules, DPPC, PSM, and Chol are depicted in Fig.
-1.
-The mole fraction of Chol in the lipid membrane was set to $x_{\mathrm{Chol}}=0$, 0.1, and 0.5, with the number of water molecules maintained at a ratio of 100 per total number of lipid and cholesterol molecules (see Table 1).
-The CHARMM36 force field was employed for the lipid molecules (DPPC, PSM, and Chol), [58] and water was modeled using the TIP3P model optimized for compatibility with the CHARMM force field.
-[59] All the MD simulations were performed using GROMACS 2022.4.
-[60] The initial configurations of the lipid membrane systems were constructed using the CHARMM-GUI.
-[61, 62, 63, 64] System equilibration was carried out following the CHARMM-GUI protocol, as described in our previous study.
-[55] Three initial configurations were prepared for each $x_{\mathrm{Chol}}$.
-The temperatures investigated were 303 K and 323 K.
-The production runs of 10 ns in the $NPT$ ensemble.
-Each production run was performed three times per configuration with different initial atomic velocities assigned according to the Maxwell–Boltzmann distribution, resulting in a total of nine trajectories for each combination of $x_{\mathrm{Chol}}$ and temperature.
+[…]
+The MD setting are the same as that in our previous study.
+The time step was set to 2 fs.
+The system temperature was controlled using the Nosé–Hoover thermostat with a coupling constant of 1 ps.
+[68, 69] We employed Parrinello–Rahman barostat with a coupling constant of 5 ps and a compressibility of $4.5\times 10^{-5}~\text{bar}^{-1}$[70] Long-range electrostatic interactions were treated using the smooth Particle-Mesh Ewald (SPME) method with a 1.2 nm cut-off.
+[71] Short-range van der Waals interactions were handled with a force-switch modifier between 1.0 and 1.2 nm.
+Neighbor lists were updated every 20 steps using the Verlet cutoff scheme.
+All bonds involving hydrogen atoms were constrained using the LINCS algorithm to allow the 2 fs time step.
+[72]
 III Results and discussion
@@
 Thus, increasing $x_{\mathrm{Chol}}$ led to a corresponding increase in the simulation box size.
-It is also observed that $x_{\mathrm{Chol}}$ influences the bending behavior of membranes and the orientation of the hydrophobic group of the lipid molecules.
-The top-down snapshots in Fig.
+As shown in Fig.
+2(a), the pure DPPC membrane at 303 K exhibits the periodic modulation of the bilayer surface, a characteristic feature of ripple phase.
+The phase behavior we obtained is consistent with the experimental findings, as 303 K is close to the temperature range of 308–314 K, in which the ripple phase is experimentally observed.
+[74, 75] The top-down snapshots in Fig.
 2 reveal that at $x_{\mathrm{Chol}}=0.1$, Chol is stably incorporated into the membrane and exhibits strong lateral aggregation, with minimal exposure at the membrane interface.
 In contrast, at $x_{\mathrm{Chol}}=0.5$, the presence of Chol becomes apparent at the membrane interface, which is regarded as the effective dilution of lipid molecules compared to the system with $x_{\mathrm{Chol}}=0.1$.
+From the radial distribution functions (RDFs) on the $xy$-plane between the phosphorus atoms and between phosphorus and cholesterol oxygen ($\mathrm{O}^{\mathrm{Chol}}$) atoms, we confirm that no macroscopic phase separation is observed and that the lipids (DPPC and PSM) and cholesterol are miscible at $x_{\mathrm{Chol}}=0$, 0.1, and 0.5, as the RDFs converge to unity in the long-range limit for all cases (see Figs.
+S2 and S3 of the supplementary material).
 We examined how the membrane area in the $xy$ plane varies with the Chol concentration.
@@
 The results of $x_{\mathrm{Chol}}=0$ and 0.1 are consistent with those reported in our previous study.
-[55] Figure S1 of the supplementary material shows the results at 323 K.
+[56] Figure S4(a) and (b) of the supplementary material shows the results at 323 K.
 Here, the $z^{\prime}$ axis is divided into three distinct regions: the membrane interior (Region 1), the interface (Region 2), and the bulk aqueous phase (Region 3).
-[31, 33] We defined Region 1 as $z^{\prime}<-0.25$ nm, Region 2 as $-0.25~\mathrm{nm}<z^{\prime}<0.85\mathrm{nm}$, and Region 3 as $z^{\prime}>0.85~\mathrm{nm}$, as indicated by the dashed vertical lines in Fig.
+[32, 34] We defined Region 1 as $z^{\prime}<-0.25$ nm, Region 2 as $0.25~\mathrm{nm}<z^{\prime}<0.85\mathrm{nm}$, and Region 3 as $z^{\prime}>0.85~\mathrm{nm}$, as indicated by the dashed vertical lines in Fig.
 4.
 Both DPPC and PSM exhibit a non-monotonic dependence on $x_{\mathrm{Chol}}$ in Region 2 at 303 K.
-In particular, for DPPC, the number density of water at the membrane interface decreased from $x_{\mathrm{Chol}}=0$ to $x_{\mathrm{Chol}}=0.1$, but increases markedly at $x_{\mathrm{Chol}}=0.5$.
-This increase in $\rho(z^{\prime})$ is consistent with the results performed by Elola et al.
-[49] As illustrated in Fig.
-S1 of the supplementary material, at 323 K, the number density of water in Region 2 shows little variation between $x_{\mathrm{Chol}}=0$ and $x_{\mathrm{Chol}}=0.1$.
+In particular, for DPPC, the number density of water at the membrane interface decreases from $x_{\mathrm{Chol}}=0$ to $x_{\mathrm{Chol}}=0.1$, but increases markedly at $x_{\mathrm{Chol}}=0.5$.
+This increase in $\rho(z^{\prime})$ is consistent with the results performed by Elola and Rodriguez [50] As illustrated in Fig.
+S4 of the supplementary material, at 323 K, the number density of water in Region 2 shows little variation between $x_{\mathrm{Chol}}=0$ and $x_{\mathrm{Chol}}=0.1$.
 However, at $x_{\mathrm{Chol}}=0.5$, it increases noticeably, similar to the trend observed at 303 K.
@@
 Consequently, the exclusion effect of Chol on water molecules is less pronounced in PSM.
-Eventually, as the Chol concentration increases, the lateral spacing between lipid headgroups expands, promoting the redistribution of water molecules from the membrane interior toward the interface.
-We next examined the H-bond number, $N_{\mathrm{HB}}$, formed by water molecules, as a function of $z^{\prime}$, as is shown in Fig.
-4 (c) and (d), for DPPC and PSM, respectively.
-The corresponding results at 323 K are displayed in Fig.
-S1 (c) and (d) of the supplementary material.
-An H-bond was identified based on the geometric criteria: an oxygen-oxygen distance $r_{\mathrm{oo}}<0.35~\mathrm{nm}$ and a hydrogen-donor-acceptor angle $0^{\circ}<\beta<30^{\circ}$[68] Note that when counting $N_{\mathrm{HB}}$, we included not only those between water molecules but also those formed between water molecules and lipid oxygen atoms, ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D6}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, as well as Chol oxygen ($\mathrm{O}^{\mathrm{Chol}}$) (see Fig.
-1).
-The value of $N_{\mathrm{HB}}$ decreases progressively towards the membrane interior in both DPPC and PSM.
-In Region 1, the presence of Chol noticeably reduces $N_{\mathrm{HB}}$, compared to the Chol-free system.
-Figure 5: Two-dimensional plots of joint distribution of $\rho(z^{\prime},\cos\theta)$, normalized by $\rho(z^{\prime})$, at 303K for $x_{\mathrm{Chol}}=0$ [(a) and (d)], 0.1 [(b) and (e)], and 0.5 [(c) and (f)].
+As the Chol concentration increases up to $x_{\mathrm{Chol}}=0.5$, the lateral spacing between lipid headgroups expands, leading to a shift of the rising edge of the water density profile toward the interfacial region.
+[…]
+Figure 5: Left panel: Schematic illustration of water orientation.
+Right panels: Two-dimensional plots of joint distribution of $\rho(z^{\prime},\cos\theta)$, normalized by $\rho(z^{\prime})$, at 303K for $x_{\mathrm{Chol}}=0$ [(a) and (d)], 0.1 [(b) and (e)], and 0.5 [(c) and (f)].
 Panels (a), (b), and (c) correspond to DPPC, while panels (d), (e), and (f) corresponds to PSM.
@@
 Specifically, the number of oxygen atoms located within $r_{\mathrm{OO}}<0.35$ nm is denoted as $N_{\mathrm{NN}}$, representing the number of nearest neighbors.
-The normalized number of H-bonds, $N_{\mathrm{HB}}/N_{\mathrm{NN}}$, was then evaluated as a function of $z^{\prime}$, with the results shown in Fig.
+The normalized number of H-bonds, $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$, was then evaluated as a function of $z^{\prime}$, with the results shown in Fig.
 4(e) and (f).
-Note that $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ denotes the average fraction of neighboring oxygen atoms that form H-bonds at each position $z^{\prime}$[69, 70] In the bulk region (Region 3), the value of $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ is approximately 0.6, indicating that about 60% of the oxygen atoms in the nearest neighbors participate in H-bonding.
-The value of $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ increases toward the interior of the membrane (Region 1) in the Chol-free system.
+The corresponding results at 323 K are displayed in Fig.
+S4 (e) and (f) of the supplementary material.
+Note that $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ denotes the average fraction of neighboring oxygen atoms that form H-bonds at each position $z^{\prime}$[82, 83] In the bulk region (Region 3), the value of $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ is approximately 0.6, indicating that about 60% of the oxygen atoms in the nearest neighbors participate in H-bonding.
+The value of $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ increases toward the interior of the membrane (Region 1) in the Chol-free system.
 In other words, when oxygen atoms are present nearby, H-bonds are formed with a higher probability than in the bulk, primarily due to the proximity of oxygen atoms from lipid molecules.
-However, at $x_{\mathrm{Chol}}=0.5$, the Chol suppressed H-bond formation in the membrane interior for both DPPC and PSM.
-This suppression can be attributed to the reduced number of H-bonds in that region caused by Chol, as shown in Fig.
-4(c) and (d).
+However, at $x_{\mathrm{Chol}}=0.5$, $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ is reduced in the membrane interior, even though $\Braket{N_{\mathrm{HB}}}$ is hardly changed.
+Since $\Braket{N_{\mathrm{NN}}}$ is higher at $x_{\mathrm{Chol}}=0.5$ than at $x_{\mathrm{Chol}}=0.1$ (see Fig.
+S5 of the supplementary material), the decrease in $\Braket{N_{\mathrm{HB}}/N_{\mathrm{NN}}}$ indicates that increasing Chol to $x_{\mathrm{Chol}}=0.5$ raises the number of neighboring oxygen atoms that do not participate in H-bonding.
 III.3 Water orientation near membrane interface
@@
 We further analyze how the orientation of water molecules at the membrane interface depends on Chol concentration.
-Specifically, we calculated the angle $\theta$ between the water dipole vector and the bilayer normal.
+Let us introduce water dipole vector $\bm{\mu}$ and the unit vector direced from membrane interface to water phase $\bm{e}_{z}$.
+We also define the azimuth angle $\theta$ between $\bm{e}_{z}$ and $\bm{\mu}$ given as $\cos\theta=\bm{e}_{z}\cdot\bm{\mu}$.
+Note that $\cos\theta>0$ and $\cos\theta\leq 0$ indicate that a water molecule is oriented toward the water phase and toward the membrane center, respectively.
 Previous MD simulations have demonstrated that water orientation varies with the nature of the interacting functional groups.
@@
 Figure 5 demonstrates that in Region 3 (bulk region) defined by $z^{\prime}>0.85$ nm, $\rho(z^{\prime},\cos\theta)$ is approximately 0.5 regardless of $\cos\theta$, indicating an isotropic orientation of water molecules.
-Furthermore, it is demonstrated that both DPPC and PSM exhibit higher values at $-1<\cos\theta<0.5$ in Region 2 defined by $-0.25~\mathrm{nm}<z^{\prime}<0.85~\mathrm{nm}$.
+Furthermore, it is demonstrated that both DPPC and PSM exhibit higher values at $1<\cos\theta<0.5$ in Region 2 defined by $0.25~\mathrm{nm}<z^{\prime}<0.85~\mathrm{nm}$.
 This Region 2 likely represents water molecules located near the phosphate groups of the lipid headgroups, as discussed in Sec III.2.
-The predominantly negative values of $\cos\theta$ suggest that the dipole vectors of these water molecules are oriented towards the membrane interior, consistent with experimental observations.
-[23] In contrast, within Region 1 defined by $z^{\prime}<-0.25$ nm, water molecules tend to orient upward, and this tendency becomes more pronounced with increasing $x_{\mathrm{Chol}}$.
+The predominantly negative values of $\cos\theta$ suggest that the dipole vectors of these water molecules are oriented toward the membrane center, consistent with experimental observations.
+[24] In contrast, within Region 1 defined by $z^{\prime}<-0.25$ nm, water molecules tend to orient toward the water phase, and this tendency becomes more pronounced with increasing $x_{\mathrm{Chol}}$.
 To highlight how their dipole alignment varies with $x_{\mathrm{Chol}}$, Fig.
@@
 The results at 323 K are illustrated in Fig.
-S3 of the supplementary material.
-As mentioned above, in Region 2, water molecules tend to orient their H-atoms toward the lipid membrane; however, the tendency weakens for $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$.
+S7 of the supplementary material.
+As mentioned above, in Region 2, water molecules tend to orient their H-atoms along the membrane normal toward the membrane center, as illustrated in Fig.
+5.
+However, this orientational tendency becomes weaker at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$.
 This reduction is likely due to the increase in the lateral spacing between lipid headgroups, as discussed in Sec.
 III.2.
-Interestingly, at $x_{\mathrm{Chol}}=0.1$, DPPC exhibited a slight increase in water molecule orientation toward the membrane interior, attributed to the condensation effect of Chol, whereas PSM showed almost no change compared to the $x_{\mathrm{Chol}}=0$ case.
+Interestingly, at $x_{\mathrm{Chol}}=0.1$, DPPC exhibited a slight increase in water molecule orientation toward the membrane center, attributed to the condensation effect of Chol, whereas PSM showed almost no change compared to the $x_{\mathrm{Chol}}=0$ case.
 These results are consistent with the observed $\rho(z^{\prime})$ dependence on $x_{\mathrm{Chol}}$ [see Fig.
 4(a) and (b)].
-In contrast, within the interior region (Region 1) of the membrane, water molecules tend to orient upward toward the membrane, particularly at $x_{\mathrm{Chol}}=0.5$.
+In contrast, within the interior region (Region 1) of the membrane, water molecules tend to orient toward the water phase, particularly at $x_{\mathrm{Chol}}=0.5$.
 This behavior can be attributed to the increased probability of H-bonding, with Chol’s hydroxyl group [denoted $\mathrm{O}^{\mathrm{Chol}}$ in Fig.
@@
 As examined in Ref.
-55, we calculated $C_{i,j}(t)$ $(i\neq j)$, the conditional probability that a water molecule initially located in Region $i$ at time 0 reaches Region $j$ by time $t$, having crossed exactly one regional boundary during this interval.
+56, we calculated $C_{i,j}(t)$ $(i\neq j)$, the conditional probability that a water molecule initially located in Region $i$ at time 0 reaches Region $j$ by time $t$, having crossed exactly one regional boundary during this interval.
 $C_{i,i}(t)$ represents the probability that the water molecule remains in Region $i$ without entering the other regions during the time interval between 0 and $t$.
@@
 The corresponding results at 323 K are shown in Fig.
-S4 of the supplementary material.
+S8 of the supplementary material.
 When $x_{\mathrm{Chol}}$ increases from 0 to 0.1, the water dynamics slowed down, indicating that the time required to transition into Region 2 became longer.
 This result is consistent with the finding reported in Ref.
-55.
+56.
 In contrast, at $x_{\mathrm{Chol}}=0.5$, the transition from Region 1 to Region 2 was accelerated for both DPPC and PSM.
@@
 The corresponding results at 323 K are shown in Fig.
-S5 of the supplementary material.
+S9 of the supplementary material.
 As time $t$ increases, the value of $C_{2,1}(t)$ saturated in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5 for both DPPC and PSM.
-Specifically, at $x_{\mathrm{Chol}}=0.5$, $C_{2,1}(t)$ converges to approximately 0.21, which is about twice the value observed at $x_{\mathrm{Chol}}=0$ This result corresponds to $C_{1,2}(t)$ shown in Fig.
+Specifically, at $x_{\mathrm{Chol}}=0.5$, $C_{2,1}(t)$ converges to approximately 0.21, which is about twice the value observed at $x_{\mathrm{Chol}}=0.1$.
+This result corresponds to $C_{1,2}(t)$ shown in Fig.
 7, indicating that the exchange between the membrane interior region (Region 1) and the interface region (Region 2) becomes faster with increasing $x_{\mathrm{Chol}}$.
@@
 These results of $C_{i,j}(t)$ indicate that the hydration state in the interface region (Region 2) is significantly influenced by Chol concentration.
-In particular, the findings suggest that water molecules in Region 2 become more stable at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$ and 0.1.
+Specifically, the findings suggest that water molecules in Region 2 become more stable at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$ and 0.1.
 III.5 H-bond lifetime between water molecules and functional groups
-Figure 9: Ratio of H-bond lifetimes, $\tau_{\mathrm{HB}}$, at $x_{\mathrm{Chol}}=0.1$ and 0.50.50.5 to those at $x_{\mathrm{Chol}}=0$, denoted as $\tau_{\mathrm{HB(Chol)}}/\tau_{\mathrm{HB(pure)}}$, for (a) DPPC and (b) PSM.
+Figure 9: Ratio of H-bond lifetimes, $\tau_{\mathrm{HB}}$, at $x_{\mathrm{Chol}}=0.1$ and $0.5$ to those at $x_{\mathrm{Chol}}=0$, denoted as $\tau_{\mathrm{HB(Chol)}}/\tau_{\mathrm{HB(pure)}}$, for (a) DPPC and (b) PSM.
 Finally, to elucidate the interfacial dynamics at a more microscopic level, we analyzed the lifetime of H-bonds formed between water molecules as well as between water molecules and functional groups of DPPC and PSM.
@@
 III.2.
-The acceptor oxygen of water molecule is denoted as $\mathrm{O}^{\mathrm{w}}$.
-Note that the donor oxygen atoms are limited to ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, respectively.
+Note that the acceptor oxygen atoms are limited to ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, respectively.
+We also note that, in the case of the H-bond between water and the nitrogen atom in PSM, water acts only as an H-bond acceptor (see Fig.
+S10 of the supplementary material for the 2D PMF using $r_{\mathrm{OO}}$ and $\beta$).
 This selection is based on the significantly lower number of H-bonds involving other oxygen atoms such as ($\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D6}}$) for DPPC and $\mathrm{O}^{\mathrm{2}}$ for PSM, as confirmed from the joint probability distribution of $r_{\mathrm{OO}}$ and $\beta$ reported in Ref.
-55.
+56.
 We define the H-bond index $h_{i,j}(t)$ is defined as 1 if atoms $i$ and $j$ form a H-bond at time $t$, and 0 otherwise, Using this definition, we calculated the H-bond time correlation function as follows:
 $P_{\mathrm{HB}}(t)=\dfrac{\langle h_{i,j}(t)h_{i,j}(0)\rangle}{\langle h_{i,j}(0)\rangle},$ (1)
-[…]
+where $\langle\cdots\rangle$ represents the ensemble average over all pairs $(i,j)$ and initial times.
+[90, 78, 91] A similar H-bond dynamics analysis was previously reported in polymer–water mixtures.
+[92]
 We computed $P_{\mathrm{HB}}(t)$ using the Monte Carlo bootstrap method from independent nine trajectories.
 [93] From the fitting of $P_{\mathrm{HB}}(t)$ to the Kohlrausch–Williams–Watts function, $P_{\mathrm{HB}}(t)=\exp[-(t/\tau_{\mathrm{KWW}})^{\beta_{\mathrm{KWW}}}]$, the H-bond lifetime $\tau_{\mathrm{HB}}$ was evaluated as
-[…]
+$\tau_{\mathrm{HB}}=\int^{\infty}_{0}P_{\mathrm{HB}}(t)dt=\dfrac{\tau_{\mathrm{KWW}}}{\beta_{\mathrm{KWW}}}\Gamma\left(\dfrac{1}{\beta_{\mathrm{KWW}}}\right),$ (2)
 where $\Gamma(\cdots)$ represents the Gamma function.
@@
 A similar tendency is observed for PSM, with a more pronounced reduction in $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.5$, which exhibits comparable values between 303 K and 323 K.
-This acceleration effect of Chol on $\tau_{\mathrm{HB}}$ can be attributed to the increased water content in the membrane interface (Region 2) at $x_{\mathrm{Chol}}=0.5$, as discussed previously [see Fig.
+Elola and Rodriguez reported that $\tau_{\mathrm{HB}}$ is reduced at $x_{\mathrm{Chol}}=0.3$ compared with $x_{\mathrm{Chol}}=0$ for the DPPC systems at 323 K.[50] Thus, the turning point at which the Chol effect on the H-bond dynamics shifts from deceleration to acceleration would lie between $x_{\mathrm{Chol}}=0.1$ and 0.3.
+The acceleration effect of Chol on $\tau_{\mathrm{HB}}$ can be attributed to the increased water content in the membrane interface (Region 2) at $x_{\mathrm{Chol}}=0.5$, as discussed previously [see Fig.
 4(a) and (b)].
@@
 In other words, although Chol enhances packing and can trap water within the membrane at low concentrations, at higher concentrations, its diluting effect dominates, facilitating water penetration at the interface.
-Correspondingly, the dynamics of interfacial water characterized by the transition behavior between the three regions and the H-bond lifetime showed the decrease at $x_{\mathrm{Chol}}=0.1$, followed by an acceleration at $x_{\mathrm{Chol}}=0.5$, consistent with previous MD simulations [49, 53] and spectroscopic experiments [13, 22, 23].
+Correspondingly, the dynamics of interfacial water characterized by the transition behavior between the three regions.
+H-bond lifetime increase at $x_{\mathrm{Chol}}=0.5$, consistent with previous MD simulations [50, 54] and spectroscopic experiments [14, 23, 24].
+Furthermore, given that the H-bond dynamics is retarded at $x_{\mathrm{Chol}}=0.1$, we can state that the significance of cholesterol varies non-monotonically depending on its concentration.
 In summary, The non-monotonic behavior in dynamics of water molecules near lipid membrane interface arises from the interplay between membrane packing and increased hydration induced by high Chol concentrations.
 SUPPLEMENTARY MATERIAL
-The supplementary material presents $\rho(z^{\prime})/\rho_{\mathrm{bulk}}$ and the number of H-bonds, and $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ as a function of $z^{\prime}$ at 323 K (Fig.
-S1); $\rho(z^{\prime},\cos\theta)$ normalized by $\rho(z^{\prime})$ at 323K (Fig.
-S2); $\langle\cos\theta\rangle$ as a function of $z^{\prime}$ at 323 K (Fig.
-S3); $C_{1,j}(t)$ at 323 K (Fig.
-S4); and $C_{2,j}(t)$ at 323 K (Fig.
-S5).
+[…]
 Acknowledgements.